OpenStudy (anonymous):

I need help with this Initial value problem. I'm just trying to figure out what the interval is.

OpenStudy (anonymous):

\[(p^2-64)\frac{ d^4s }{ dp^4 }+p^4(\frac{ d^3x }{ dp^3 })+\frac{ 1 }{ p^2+1 }(\frac{ ds }{ dp })+s = \sin(p) \]

OpenStudy (anonymous):

\[s(7) = 1093\]

OpenStudy (anonymous):

\[\frac{ ds }{ dp } = 10\] \[\frac{ d^2s }{ dp^2 } =3\] \[\frac{ d^3s }{ dp^3 } = 3\]

OpenStudy (aaronq):

post this in the math section?

OpenStudy (anonymous):

dammit